OpenStudy (mattyboyy):

Help please?? For the upcoming holiday season, Dorothy wants to mold 20 bars of chocolate into tiny pyramids. Each bar of chocolate contains 6 cubic inches of chocolate. Dorothy plans to make chocolate pyramids with a square base of 1 square inch and a height of 2 inches. What is the largest number of pyramids she can make from the 20 bars of chocolate? A) 240 B) 180 C) 450 D) 520 E) 580

OpenStudy (mattyboyy):

I think its B

OpenStudy (mattyboyy):

Can you prove that for me please?

OpenStudy (aggie12341):

It is B because if each bar of chocolate contains 6 cubic inches of chocolate and u have 20 bars, then in all u have 120 cubic inches V = (1/3) * b^2 * h V = (1/3) * (1)^2 * 2 V = (1/3) * 2 V = (2/3) cubic inches 120/(2/3) = 180 chocolate pyramids

OpenStudy (aggie12341):

Hope that helps! :)

OpenStudy (mattyboyy):

Thanks lol